The vertical and horizontal components of the initial velocity of a projectile are 36m/s and 64m/s find the initial velocity


1
Expert's answer
2021-01-31T15:11:33-0500
"v_x=\sqrt{v_h^2+v_v^2}=\sqrt{(36\ \dfrac{m}{s})^2+(64\ \dfrac{m}{s})^2}=73.43\ \dfrac{m}{s}."